# Cluster Magnification, Root Capacity, Unique Chains and Base Change

### v2

@@ 4.1. Strong cluster magnification
The number $[F:K]=d$ is called the magnification factor. The magnification is called trivial if $d=1$ and nontrivial otherwise.
@@
Let $L_{1}$ be Galois closure of $L/K^{\prime}$. So $L_{1}\subset\tilde{L}$. Since $\tilde{L}$ and $F$ are linearly disjoint over $K$. Thus $L_{1}$ and $F$ are linearly disjoint over $K$. Hence by Lemma 2.2.2 we have
@@
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from a subextension $L/K$. From Prop. 4.1.4 (3) and (4), we get $A=G,B=R$ and $A^{\prime}=H$ with the required conditions since, $d=|R|>1,n=[G:H]>2$.
@@
    Since $1\times B$ is normal in $G^{\prime}$, we conclude that $N:=\tilde{M}^{1\times B}$ is Galois over $K$ with Galois group $A$. Let $L:=\tilde{M}^{A^{\prime}\times B}$. Hence $L/K$ has degree $n=[A:A^{\prime}]>2$.
@@
    Since $A\times 1$ is normal in $G^{\prime}$, we conclude that $F:=\tilde{M}^{A\times 1}$ is Galois over $K$ with Galois group $B$ and degree $d=|B|$. Let $\tilde{L}$ be Galois closure of $L$ in $\bar{K}$. Since, $L\subset N$, we have $\tilde{L}\subset N$. The intersection of fields $N\cap F$ corresponds to the subgroup generated by $A\times 1$ and $1\times B$ which is $G^{\prime}$. Hence $N\cap F=K$. Thus $\tilde{L}\cap F=K$. So $\tilde{L}$ and $F$ are linearly disjoint over $K$.
@@
In the above proof of the converse part, we can additionally conclude $N=\tilde{L}$. Since, $N\cap F=\tilde{L}\cap F=K$ we get $[F:K]=[NF:N]=[\tilde{L}F:F]$. Now, $NF\subset\tilde{M}$ corresponds to intersection of the groups $A\times 1$ and $1\times B$ which is trivial. Hence $NF=\tilde{M}$. From Prop 4.1.4 (2), $\tilde{M}=\tilde{L}F$. Hence, $[N:K]=[\tilde{L}:K]$. Thus, $N=\tilde{L}$.
@@
Since $M/K$ is Galois, $\tilde{M}=M$, $G^{\prime}={\rm Gal}(M/K)$ and $H^{\prime}$ is trivial. So $A^{\prime}$ is trivial and $[A:A^{\prime}]=|A|$. Also $\tilde{L}=N=L$. Hence, $L/K$ is Galois. ∎
@@
Alternatively, we can use the criterion in Thm 4.1.5. The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures. Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.
@@ 4.2. Strong cluster magnification problem for irreducible polynomials
    By Prop. 4.1.8, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
@@
By Thm. 4.1.5, we get an equivalent criterion for strong cluster magnification of an irreducible polynomial.
@@ 4.3. Weak cluster magnification
From Def 4.1.1 and Thm 2.2.3, if $M/K$ is obtained by strong cluster magnification from $L/K$ then $M/K$ is obtained by weak cluster magnification from $L/K$. This justifies the word ‘weak’ in above definition.
@@
Let $L/K$ be a finite extension that is not Galois. So $r_{K}(L)\neq[L:K]$. We also have $r_{K}(\tilde{L})=[\tilde{L}:K]$. Now $r_{K}(L)~{}|~{}[L:K]$ and $[L:K]~{}|~{}[\tilde{L}:K]$. Hence $r_{K}(L)~{}|~{}r_{K}(\tilde{L})$ and $r_{K}(L)\neq r_{K}(\tilde{L})$. Thus $\tilde{L}/K$ is obtained by nontrivial weak cluster magnification of $L/K$.
We claim that $\tilde{L}/K$ is not obtained by strong cluster magnification of $L/K$. Assume the contrary. Then for $M=\tilde{L}$ we must have by Def. 4.1.1 and Lem. 2.2.2 that $M\cap\tilde{L}=L$ which is a contradiction since $L/K$ is not Galois. Moreover $\tilde{L}/K$ is not obtained by strong cluster magnification of $N/K$ for any $\tilde{L}\supsetneq N\supset L$.
@@ 5.1. Cluster tower of a polynomial
In [5], the notion of degree sequence and length of such tower are defined as follows.
@@
Suppose $l$ and $n$ are positive integers such that $l|n$. Consider their prime factorisations $n=\prod\limits_{p}p^{v_{p}(n)}$ and $l=\prod\limits_{p}p^{v_{p}(l)}$ with $v_{p}(l)\leq v_{p}(n)$ for every prime $p$ (here $v_{p}$ is usual p-adic valuation). Then
@@
    $\phi(n)/\phi(l)=(n/l).(\phi(k)/k)$ where $k=\prod\limits_{p\nmid l}p^{v_{p}(n)}$.
@@
    $\phi(n)/\phi(l)=n/l$ if and only if $n$ and $l$ have same prime factors.
@@
Further, assume that $n$ is composite and $l|n$, where for the cases $n$ odd or $n\equiv 0\ (\text{mod}\ 4)$, we assume $2<l<n$; and for the case $n\equiv 2\ (\text{mod}\ 4)$ we assume $22<l<n/$ 2. Because of our assumptions, $ab^{n/l}\neq\pm a,\pm ab$ and $1<\phi(l)<\phi(n)$ by (3) of Prop 5.1.2. Also, $\phi(l)|\phi(n)$ by (2) of Prop 5.1.2. Since $\mathbb{Q}(b)/\mathbb{Q}$ is Galois, by Thm. 2.6 in [2], we get
${{\rm Gal}}(\mathbb{Q}(a,b^{n/l})/\mathbb{Q}(a))\cong{{\rm Gal}}(\mathbb{Q}(b^{n/l})/\mathbb{Q})\cong(\mathbb{Z}/l\mathbb{Z})^{\times}.$
@@
Because of our assumptions, we have $(n/l)+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
$\beta_{3}=ab^{n/l},\beta_{(n/l)+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
@@ 5.2. Group theoretic formulation of cluster towers
Let $m_{1}<m_{2}<\dots<m_{l}$ be all the indices $i$ such that $J_{i}\neq J_{i-1}$. The length of above cluster tower is $l+2$. Here, $m_{l}$ is smallest index $i$ such that $K_{i}=K_{f}$.
@@ 6.1. The group of automorphisms of finite extensions
Suppose $M/L$ is Galois. Then we can replace $M^{{\rm Aut}(M/L)}$ by $L$ in above Prop. 6.1.3. If additionally we have ${\rm{\rm Aut}}(M/L)\unlhd\ {\rm Aut}(M/K)$, then $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. This is converse to Prop 6.1.1 (2) under the condition $M/L$ is Galois.
@@ 6.2. Root capacity
Let $L/K$ be an extension. By primitive element theorem $L=K(\alpha)$ for some $\alpha\in\bar{K}$. We define $\rho_{K}(M,L):=\rho_{K}(M,\alpha)$ which is well defined by the following proposition. We call this quantity as the root capacity of $M$ over $L$ (with base field $K$ fixed).
@@
Let $\alpha$ and $f$ be as in Def. 6.2.1. Now, $\rho_{K}(K(\alpha),\alpha)=r_{K}(f)$. Thus the above proposition proves that given an extension $L/K$, the cluster size is the same for all irreducible polynomials that are the minimal polynomials of primitive elements of $L$ over $K$ (Corollary 1 in [5]). This proof is independent of Thm 2.1.1 (3).
@@
Let $L_{k}$ be as in proof of Thm 3.2.4. Let $L=L_{1}$. So $r_{K}(L)=1,s_{K}(L)=n$ and $\tilde{L}=L_{n-1}$. Also $\rho_{K}(L_{k},L)=k$ for $1\leq k\leq(n-2)$. Also $\rho_{K}(L_{2},L)=r_{K}(L_{2})=2$.
@@
Consider the set of $s$ many representatives of clusters of minimal polynomial of a primitive element of $L/K$. Consider the subsets of representatives contained in $M_{1}$ and $M_{2}$ and let $a$ and $b$ be their respective cardinalities. Hence $\rho_{K}(M_{1})=a.r_{K}(L)$ and $\rho_{K}(M_{2})=b.r_{K}(L)$. Let $i$ be cardinality of intersection of these two sets. Hence $\rho_{K}(M_{1}\cap M_{2},L)\geq i.r_{K}(L)$. Also since cardinality of union of these two sets is $a+b-i$. Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$. Thus
@@
    $M\cap\tilde{L}=L\implies\rho_{K}(M,L)=r_{K}(L)$.
@@
Proof of (7): If the extension $M/K$ is obtained by strong cluster magnification from $L/K$, then we have $M\cap\tilde{L}=L$ (see Def.  4.1.1 and Lem. 2.2.2 ).
@@
Suppose $M\cap\tilde{L}=L$. Hence from Prop 6.2.9 (6), Prop 6.2.4 (2) and Prop 6.1.1 (2), we have
@@
Remark 6.2.11.
The converse of above statement is not true. Take $M=\tilde{L}$ where $L/K$ is not a Galois extension. In this case, $r_{K}(L)<[L:K]$ and hence
$r_{K}(\tilde{L})/r_{K}(L)=[\tilde{L}:K]/r_{K}(L)>[\tilde{L}:K]/[L:K]=[\tilde{L}:L].$

### v3

@@ 4.1. Strong cluster magnification
The number $[F:K]$ is called the magnification factor and denoted by $d$. The magnification is called trivial if $F=K$ and nontrivial otherwise.
@@
Remark 4.1.3.
Let $LF/K$ be obtained by strong cluster magnification from $L/K$ through $F/K$. If $K\subset L^{\prime}\subset L$. Then $L^{\prime}F/K$ is obtained by strong cluster magnification from $L^{\prime}/K$ through $F/K$.
@@
Let $L_{1}$ be Galois closure of $L/K^{\prime}$. So $L_{1}\subset\tilde{L}$. Since $\tilde{L}$ and $F$ are linearly disjoint over $K$, we conclude that $L_{1}$ and $F$ are linearly disjoint over $K$. Hence by Lemma 2.2.2 we have
@@
In view of property (1) in Prop. 4.1.5, we see that ${r_{K}(L)}|{r_{K}(M)}$ and the ratio ${r_{K}(M)}/{r_{K}(L)}$ is indeed same as the degree $d$ = $[F:K]$, which is the magnification factor for $M/L$ over $K$ as defined earlier.
@@
Suppose $M/K$ is obtained by nontrivial strong cluster magnification from a subextension $L/K$. From Prop. 4.1.5 (3) and (4), we get $A=G,B=R$ and $A^{\prime}=H$ with the required conditions since, $d=|R|>1,n=[G:H]>2$.
@@
    Since $1\times B$ is normal in $G^{\prime}$, we conclude that $\tilde{M}_{B}:=\tilde{M}^{1\times B}$ is Galois over $K$ with Galois group $A$. Let $L:=\tilde{M}^{A^{\prime}\times B}$. Hence $L/K$ has degree $n=[A:A^{\prime}]>2$.
@@
    Since $A\times 1$ is normal in $G^{\prime}$, we conclude that $F:=\tilde{M}^{A\times 1}$ is Galois over $K$ with Galois group $B$ and degree $d=|B|$. Let $\tilde{L}$ be Galois closure of $L$ in $\bar{K}$. Since, $L\subset\tilde{M}_{B}$, we have $\tilde{L}\subset\tilde{M}_{B}$. The intersection of fields $\tilde{M}_{B}\cap F$ corresponds to the subgroup generated by $A\times 1$ and $1\times B$ which is $G^{\prime}$. Hence $\tilde{M}_{B}\cap F=K$. Thus $\tilde{L}\cap F=K$. So $\tilde{L}$ and $F$ are linearly disjoint over $K$.
@@
In the above proof of the converse part, we can additionally conclude $\tilde{M}_{B}=\tilde{L}$. Since, $\tilde{M}_{B}\cap F=\tilde{L}\cap F=K$ we get $[F:K]=[\tilde{M}_{B}F:\tilde{M}_{B}]=[\tilde{L}F:F]$. Now, $\tilde{M}_{B}F\subset\tilde{M}$ corresponds to intersection of the groups $A\times 1$ and $1\times B$ which is trivial. Hence $\tilde{M}_{B}F=\tilde{M}$. From prop. 4.1.5 (2), $\tilde{M}=\tilde{L}F$. Hence, $[\tilde{M}_{B}:K]=[\tilde{L}:K]$. Thus, $\tilde{M}_{B}=\tilde{L}$.
@@
Since $M/K$ is Galois, $\tilde{M}=M$, $G^{\prime}={\rm Gal}(M/K)$ and $H^{\prime}$ is trivial. So $A^{\prime}$ is trivial and $[A:A^{\prime}]=|A|$. Also $\tilde{L}=\tilde{M}_{B}=L$. Hence, $L/K$ is Galois. ∎
@@
Alternatively, we can use the criterion in Thm. 4.1.6. The isomorphism between $M$ and $M^{\prime}$ extends to isomorphism between their Galois closures. Hence we also get an isomorphism between $G^{\prime},H^{\prime}$ and corresponding groups of $M^{\prime}$.
Remark 4.1.10.
From the proof of Thm. 4.1.6, we get the following way to construct all fields $M/K$ which are obtained by nontrivial strong cluster magnification from some subextension $L/K$.
Suppose $A\times B$ is realisable as a Galois group over $K$ for nontrivial groups $A$ and $B$ with a subgroup $A^{\prime}\subset A$ with $[A:A^{\prime}]>2$ such that $\bigcap\limits_{a\in A}\ aA^{\prime}a^{-1}=1$. Let $P$ be such that ${\rm Gal}(P/K)=A\times B$. Then our required fields are $M=P^{A^{\prime}\times 1}$ and $L=P^{A^{\prime}\times B}$.
@@ 4.2. Strong cluster magnification problem for irreducible polynomials
    By Prop. 4.1.9, if for some $i$, $M_{i}/K$ is obtained by strong cluster magnification from $L_{i}/K$ then for each $1\leq j\leq s$, $M_{j}/K$ is obtained by strong cluster magnification from some subextension $L_{j}/K$.
@@
By Thm. 4.1.6, we get an equivalent criterion for strong cluster magnification of an irreducible polynomial.
@@ 4.3. Weak cluster magnification
From Def 4.1.1 and Thm. 2.2.3, if $M/K$ is obtained by strong cluster magnification from $L/K$ then $M/K$ is obtained by weak cluster magnification from $L/K$. This justifies the word ‘weak’ in above definition.
@@
[…]
We claim that $M/K$ is not obtained by strong cluster magnification from $L/K$. Assume the contrary. Then we must have by Def. 4.1.1 and Lem. 2.2.2 that $M\cap\tilde{L}=L$ which is a contradiction since $L/K$ is not Galois. Moreover $M/K$ is not obtained by strong cluster magnification of $L_{1}/K$ for any $\tilde{L}\supsetneq L_{1}\supset L$.
Example 4.3.4.
Let $L_{k}$ be as in proof of Thm. 3.2.4. Then for $j<k$, $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not obtained by nontrivial strong cluster magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$.
One can also verify that $[L_{k}:L_{j}]=r_{K}(L_{k})/r_{K}(L_{j})\iff k>n/2\ \text{and}\ j=n-k$.
Example 4.3.5.
Let $K=\mathbb{Q}$ and $\xi_{n}$ be $n$-th primitive root of unity.
  1. (1)
[…]
    We can use a similar argument as above to conclude that for integers $k>j\geq 3$ for prime $p=2$ and integers $k>j\geq 2$ for prime $p\neq 2$, $M=\mathbb{Q}(\xi_{p^{k}})$ is not obtained by nontrivial strong cluster magnification from $L=\mathbb{Q}(\xi_{p^{j}})$.
  2. (2)
    Let $n,l$ be integers such that $6<l<n$, $l|n$ such that $n=lm$ and $\text{gcd}(l,m)=1$. Let $M=\mathbb{Q}(\xi_{n})$ and $L=\mathbb{Q}(\xi_{l})$. We have
    $\begin{split}{\rm Gal}(M/K)&\cong\prod\limits_{p|n}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}=\prod\limits_{p|l}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}\\ &\cong{\rm Gal}(L/K)\ \times\prod\limits_{p|m}\ (\mathbb{Z}/p^{v_{p}(n)}\mathbb{Z})^{\times}.\end{split}$
    Hence by Corollary 4.1.8, $M/K$ is obtained by nontrivial strong cluster magnification from $L/K$ through $F/K$ where $F=\mathbb{Q}(\xi_{m})=\mathbb{Q}(\xi_{n/l})$.
@@ 5.1. Cluster tower of a polynomial
In [5], the notions of degree sequence and length of such tower are defined as follows.
@@
Suppose $l$ and $n$ are positive integers such that $l|n$ with $n=lm$. Consider their prime factorisations $n=\prod\limits_{p}p^{v_{p}(n)}$ and $l=\prod\limits_{p}p^{v_{p}(l)}$ with $v_{p}(l)\leq v_{p}(n)$ for every prime $p$ (here $v_{p}$ is usual $p$-adic valuation). Then
@@
    $\phi(n)/\phi(l)=m\phi(k)/k$ where $k=\prod\limits_{p\nmid l}p^{v_{p}(n)}$.
@@
    $\phi(n)/\phi(l)=m$ if and only if $n$ and $l$ have same prime factors.
@@
Further, assume that $n$ is composite and $l|n$ with $n=ml$, where for the cases $n$ odd or $n\equiv 0\ (\text{mod}\ 4)$, we assume $2<l<n$; and for the case $n\equiv 2\ (\text{mod}\ 4)$ we assume $22<l<n/$ 2. Because of our assumptions, $ab^{m}\neq\pm a,\pm ab$ and $1<\phi(l)<\phi(n)$ by (3) of Prop. 5.1.2. Also, $\phi(l)|\phi(n)$ by (2) of Prop. 5.1.2. Since $\mathbb{Q}(b)/\mathbb{Q}$ is Galois, by Thm. 2.6 in [2], we get
${{\rm Gal}}(\mathbb{Q}(a,b^{m})/\mathbb{Q}(a))\cong{{\rm Gal}}(\mathbb{Q}(b^{m})/\mathbb{Q})\cong(\mathbb{Z}/l\mathbb{Z})^{\times}.$
@@
Because of our assumptions, we have $m+1\leq s$. Let the representatives of clusters of roots of $f$ be given by
$\beta_{3}=ab^{m},\beta_{m+1}=ab^{2},\beta_{i}=ab^{i-1}\ \text{ for other}\ 1\leq i\leq s.$
@@ 5.2. Group theoretic formulation of cluster towers
Let $m_{1}<m_{2}<\dots<m_{l}$ be all the indices $i>1$ such that $J_{i}\neq J_{i-1}$. The length of above cluster tower is $l+2$. Here, $m_{l}$ is smallest index $i$ such that $K_{i}=K_{f}$.
@@
Example 5.2.2.
Let $f$ be an irreducible polynomial such that $|G|=np$ and $|H|=p$ where $p\nmid n$. Then length of cluster tower is $3$ and degree sequence is $n,np$. Both degree sequence and length of cluster tower are independent of the ordering of the $\beta_{i}$’s. As a particular case, $G=\mathfrak{A}_{4}$ for a degree-$4$ irreducible polynomial $f$.
@@ 6.1. The group of automorphisms of finite extensions
Suppose $M/L$ is Galois. Then we can replace $M^{{\rm Aut}(M/L)}$ by $L$ in above Prop. 6.1.3. If additionally we have ${\rm{\rm Aut}}(M/L)\unlhd\ {\rm Aut}(M/K)$, then $\sigma|_{L}\in{\rm Aut}(L/K)$ for any $\sigma\in{\rm Aut}(M/K)$. This is converse to Prop. 6.1.1 (2) under the condition $M/L$ is Galois.
@@ 6.2. Root capacity
Let $L/K$ be an extension. By primitive element theorem $L=K(\alpha)$ for some $\alpha\in\bar{K}$. We define $\rho_{K}(M,L):=\rho_{K}(M,\alpha)$ which is well defined by the following proposition. We call this quantity as the root capacity of $M$ with respect to $L$ (with base field $K$ fixed).
@@
Let $\alpha$ and $f$ be as in Def. 6.2.1. Now, $\rho_{K}(K(\alpha),\alpha)=r_{K}(f)$. Thus the above proposition proves that given an extension $L/K$, the cluster size is the same for all irreducible polynomials that are the minimal polynomials of primitive elements of $L$ over $K$ (Corollary 1 in [5]). This proof is independent of Thm. 2.1.1 (3).
@@
Let $L_{k}$ be as in proof of Thm. 3.2.4. Let $L=L_{1}$. So $r_{K}(L)=1,s_{K}(L)=n$ and $\tilde{L}=L_{n-1}$. Also $\rho_{K}(L_{k},L)=k$ for $1\leq k\leq(n-2)$. Also $\rho_{K}(L_{2},L)=r_{K}(L_{2})=2$.
Example 6.2.8.
Let notation be as in subsection 5.2. Let $L=K_{1}$ and $m_{0}=1$. Then for $i\geq 0$ and $m_{i}\leq m<m_{i+1}$, we have $\rho(K_{m},L)=\rho(K_{m_{i}},L)=(m_{i+1}-1)r_{K}(L)$.
@@
Consider the set of $s$ many representatives of clusters of minimal polynomial of a primitive element of $L/K$. Consider the subsets of representatives contained in $M_{1}$ and $M_{2}$ and let $a$ and $b$ be their respective cardinalities. Hence $\rho_{K}(M_{1})=a.r_{K}(L)$ and $\rho_{K}(M_{2})=b.r_{K}(L)$. Let $i$ be cardinality of intersection of these two sets. Hence $\rho_{K}(M_{1}\cap M_{2},L)\geq i.r_{K}(L)$. Also since cardinality of union of these two sets is $a+b-i$. Hence $\rho_{K}(M_{1}M_{2},L)\geq(a+b-i).r_{K}(L)$. Hence we are done.
Example 6.2.10.
Let $p$ be a prime and let $a=\sqrt[12]{p}$ and $b=\xi_{12}$. Let $L=\mathbb{Q}(a)$ and $M_{1}=\mathbb{Q}(a,ab^{2})$ and $M_{2}=\mathbb{Q}(a,ab^{3})$. Thus $M_{1}M_{2}=\tilde{L}$. So $\rho_{K}(M_{1}M_{2},L)=12,\rho_{K}(M_{1},L)=6,\rho_{K}(M_{2},L)=4$. Hence $\rho_{K}(M_{1}M_{2},L)>\rho_{K}(M_{1},L)+\rho_{K}(M_{2},L)-\rho_{K}(M_{1}\cap M_{2},L)$. The inequality is strict here.
@@
    $M\cap\tilde{L}=L\implies\rho_{K}(M,L)=r_{K}(L)$. Also $M\cap\tilde{L}=K\implies\rho_{K}(M,L)=0$.
@@
Proof of (7): If the extension $M/K$ is obtained by strong cluster magnification from $L/K$, then we have $M\cap\tilde{L}=L$ (see Def.  4.1.1 and Lem. 2.2.2 ). ∎
Remark 6.2.12.
The statement $L_{M}=M\cap\tilde{L}$ and the statement $\rho_{K}(M,L)=0\implies M\cap\tilde{L}=K$ are not true in general. Let $K=\mathbb{Q},L=\mathbb{Q}(\sqrt[4]{2})$ and $M=\mathbb{Q}(\sqrt[6]{2})$. Hence $\tilde{L}=\mathbb{Q}(\sqrt[4]{2},\iota)$ and $M\cap\tilde{L}=\mathbb{Q}(\sqrt[]{2})$ and $L_{M}=\mathbb{Q}$.
@@
Suppose $M\cap\tilde{L}=L$. Hence from Prop. 6.2.11 (6), Prop. 6.2.4 (2) and Prop. 6.1.1 (2), we have